OpenStudy (anonymous):

A garden is 3.1 times as long as it is wide. if it has a perimeter of 93.48 feet. How wide is the garden?

OpenStudy (anonymous):

Welcome to OPEN STUDY XD

OpenStudy (anonymous):

Thank you

OpenStudy (anonymous):

http://www.math.com/tables/geometry/perimeter.htm here is the site it might help..

OpenStudy (anonymous):

I just want to know how I would set up the equation?

OpenStudy (anonymous):

Suppose its width is W.. So, it is 3.1 times as long as its width.. So, if we assume L as Length, So, W = 3.1L What is the Perimeter Of Rectangle (Garden)?? \[Perimeter = 2(Length + Width)\] \[P = 2(L + W)\] But W = 3.1L \[93.48 = 2(L + 3.1L) = 8.2L\] \[L = 11.4\] So, W = 3.1L = 3.1(11.4) = 35.34..

OpenStudy (anonymous):

\[\left\{2w+2L=\frac{9348}{100},\frac{L}{w}=\frac{31}{10}\right\}\]\[\left\{w=\frac{57}{5},L=\frac{1767}{50}\right\}\to \{w=11.4,L=35.34\} \]